User Tools

Site Tools


handover:crds81

CRDS81

Hobart 26

Module: HOB+0047, 0.0GB

  • 17:39 UT - Experiment started OK (JS)
  • 17:20 Drivepc restarted in time for experiment (JS)
  • 17:20 70K stage at 110, 40k stage at 80, pressure at 4.9 will need vacuuming in the morning (JS)
  • 07:22 UT - Jim reported that the received had been cooled. The stages are currently: 79 at 70K and 59 at 20K. (Ellen)
  • 07:25 UT - Can't connect to newsmerd on monica so I can't check the He supply pressure. (Ellen)
  • 09:39 UT - S/X receiver stages are now at 67 and 53. (Ellen)
  • 09:40 UT - Tsys values in bands 01 - 08 are approx 110. This is probably due to the current storm. (Ellen)
  • 11:30 UT - Tsys values are back to normal. (Ellen)
  • 13:36 UT - Status monitor reporting date as 1969.293.01.48 UT, schedule appears unaffected. (Bryn)

Hobart 12

Module: HOB+0075, 6264 GB

  • 17:39 UT - Experiment started OK (JS)
  • 09:40 UT - Tsys values all way too high, mostly consistently around 200. This is probably due to the current storm, although Ho system temps are only up to ~110. (Ellen)
  • 11:33 UT - X-band has cooled to normal temps but S-band is still too high. (Ellen)
  • 16:37 UT - Module ran out of space - no second module to switch to. Experiment Halted. (Bryn)

Katherine Module: NRAO-020 0 GB

  • 17:39 UT - Experiment started OK (JS)

Yarragadee

Module: USN-0049, 0.000 GB

  • 17:39 UT - Experiment started OK (JS)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/crds81.txt · Last modified: 2016/01/28 16:54 by Tiege McCarthy